OpenStudy (anonymous):

Tony went to the racetrack with $80. He won $15 on the 1st race, lost $12 on the 2nd, lost $6 on the 3rd and won $8 on the 4th. How much money did he have after the 4th race.

OpenStudy (anonymous):

when he wins add,when he loses subtract...start with 80+15....

OpenStudy (anonymous):

Yep do that way........ 80+15-12-6+8=85

OpenStudy (anonymous):

15-12-6+8 = 5 so profit is $5 so..total... 80 + 5 = $85